OpenStudy (anonymous):

find the rectangular coordinates for (-4, pi/2)

OpenStudy (anonymous):

help

OpenStudy (jdoe0001):

\(\bf \left(-4,\frac{\pi}{2}\right)\qquad \qquad x=r\cdot cos(\theta)\qquad y=r\cdot sin(\theta)\qquad thus\\ \quad \\ x=(-4)cos\left(\frac{\pi}{2}\right)\qquad \qquad y=(-4)sin\left(\frac{\pi}{2}\right)\)

OpenStudy (jdoe0001):

so get the value for \(\bf \cfrac{\pi}{2}\) from your Unit Circle, and multiply away, to get the (x, y) rectangular coordinates

OpenStudy (anonymous):

thanks

OpenStudy (jdoe0001):

yw